From the top of a cliff, Tony Stark can see a boat on the ocean. The angle of depression is 41°. If Tony Stark uses his laser range finder and determines that the boat is 1.4 kilometers away, how high is the cliff?


1
Expert's answer
2021-02-24T14:21:18-0500

Solution.

Let "h" be the height of the cliff, then

"\sin{41°}=\frac{h}{1.4},"

from here

"h=1.4\sin(41°)\approx 1.4•0.656=\\=0.9184\text{ 
kilometers}=
918.4\text{ meters.}
\newline"

Answer. "0.9184\text{ 
kilometers, or }
918.4\text{ meters.}
\newline"